# Vanishing criteria for Ceresa cycles

### v1

@@ Abstract.
Let $C$ be a smooth projective curve, and let $J$ be its Jacobian. We prove vanishing criteria for the Ceresa cycle $\kappa(C)\in\mathrm{CH}_{1}(J)\otimes{\rm\mathbb{Q}}$ in the Chow group of 1-cycles on $J$. Namely,
@@ 1. $(A)$
    If $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$, then $\kappa(C)$ vanishes;
@@ 2. $(B)$
    If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then $\kappa(C)$ vanishes modulo algebraic equivalence,
@@
We then study the first interesting case where $(B)$ holds but $(A)$ does not, namely the case of Picard curves $C\colon y^{3}=x^{4}+ax^{2}+bx+c$. Using work of Schoen on the Hodge conjecture, we show that the Ceresa cycle of a Picard curve is torsion in the Griffiths group. Moreover, we determine exactly when it is torsion in the Chow group. As a byproduct, we show that there are infinitely many plane quartic curves over ${\rm\mathbb{Q}}$ with torsion Ceresa cycle (in fact, there is a one parameter family of such curves). Finally, we determine which automorphism group strata are contained in the vanishing locus of the universal Ceresa cycle over $\mathcal{M}_{3}$.
@@ 1. Introduction
in the Chow group modulo rational equivalence. If $\kappa_{C,e}$ is torsion, then $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$, where $K_{C}$ is the canonical divisor class. Moreover if $(2g-2)e$ is canonical, then the image of $\kappa_{C,e}$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$ is independent of $e$ and we denote it by $\kappa(C)$, see §2.7 for these claims. The class $\kappa(C)$ vanishes if and only if $\kappa_{C,e}$ is torsion for some degree-$1$ divisor $e$.
@@
The vanishing of the Ceresa cycle is interesting for various reasons. For example, $\kappa(C)=0$ if and only if the Chow motive $\mathfrak{h}(C)$ has a multiplicative Chow-Künneth decomposition (by [19, Proposition 3.1] and Proposition 2.12). Moreover $\bar{\kappa}(C)=0$ if and only if the tautological subring modulo algebraic equivalence is generated by a theta divisor (by [6, Corollary 3.4]), in which case Poincaré’s formula $[C]=\frac{\Theta^{g-1}}{(g-1)!}$ holds modulo algebraic equivalence. More generally, the Ceresa cycle over $\mathcal{M}_{g}$ serves as a testing ground for the study of homologically trivial algebraic cycles in codimension greater than 1.
@@ 1.1. Vanishing criteria
This improves on a recent result of Qiu and Zhang stating that if $(\operatorname{H}^{1}(C)^{\otimes 3})^{{\rm Aut}(C)}=0$, then $\kappa(C)=0$ [39].111 This formulation is equivalent to theirs, using [46, Theorem 1.5.5]. By contrast, Theorem A requires only the weaker condition that the subrepresentation $\operatorname{H}^{3}(J)_{\mathrm{prim}}\subset\operatorname{H}^{3}(J)\simeq\bigwedge^{3}\operatorname{H}^{1}(C)\subset\operatorname{H}^{1}(C)^{\otimes 3}$ has no nontrivial ${\rm Aut}(C)$-fixed points. If the quotient $C/{\rm Aut}(C)$ has genus 00, then our hypothesis is equivalent to $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$.
@@
Our second result is a vanishing criterion for $\overline{\kappa}(C)$, however our proof works only in characteristic 00 and is conditional on the Hodge conjecture:
@@
Suppose $\mathrm{char}(k)=0$ and assume the Hodge conjecture for abelian varieties. If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then $\bar{\kappa}(C)=0$.
More precisely, we require the Hodge conjecture for $J\times A$, for a specific abelian variety $A$ described in Proposition 3.3. Note that $\operatorname{H}^{0}(J,\Omega_{J}^{3})\simeq\bigwedge^{3}\operatorname{H}^{0}(C,\Omega_{C}^{1})$, so the conditions of Theorems A (when $\mathrm{char}(k)=0$) and B both depend only on the abstract representation $(G,V)=({\rm Aut}(C),\operatorname{H}^{0}(C,\Omega_{C}^{1}))$.
@@ 1.2. Picard curves
The condition $\operatorname{H}^{3}(J)_{\mathrm{prim}}^{{\rm Aut}(C)}=0$ in Theorem A is only rarely satisfied, e.g. it holds for exactly two plane quartic curves over ${\rm\mathbb{C}}$ (see Theorem D below). However, the condition $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$ of Theorem B is satisfied for all Picard curves $y^{3}=x^{4}+ax^{2}+bx+c$. For such curves it turns out that we require a nontrivial case of the Hodge conjecture to apply Theorem B, namely that a certain Weil class on an abelian fourfold is algebraic. Fortunately, Schoen has proved the Hodge conjecture in our specific situation [41]. (See also recent work of Markman [32].) This leads to an unconditional proof of the vanishing of $\bar{\kappa}(C)$. By further analyzing $\kappa(C)$ we prove the following precise characterization of the vanishing of $\kappa(C)$ in the Picard family:
@@
Suppose $\mathrm{char}(k)=0$ and let $C_{f}$ be a smooth projective Picard curve with model
@@
Here, ${\rm disc}(f)$ is the discriminant of $f$ and the coordinates of $P_{f}$ are the usual $I$- and $J$-invariants of the binary quartic form $z^{4}f(x/z)$; see §4.1. The fact that $P_{f}$ defines a point on $E_{f}$ follows from a classical relation between the invariants of a binary quartic. See also §4.2 for a more precise version of Theorem C, which relates order of $\kappa_{C_{f},P_{\infty}}$ to the order of $P_{f}$, and in particular shows that $\kappa_{C_{f},P_{\infty}}$ can have arbitrarily large (finite) order in this family. Theorem C also gives one-parameter families of plane quartic curves with vanishing $\kappa(C)$, corresponding to components of the torsion locus of $P_{f}$. Such families exist even over ${\rm\mathbb{Q}}$, e.g. the family $y^{3}=x^{4}-12x^{2}+tx-12$.
Theorem C generalizes the results of [27] concerning bielliptic Picard curves, i.e. those with $b=0$ in (1.2). There we exploited the explicit geometry of the bielliptic cover to characterize the vanishing locus of $\kappa(C)$. Here we must proceed more indirectly: using Schoen’s work on the Hodge conjecture, we first show that the vanishing of $\kappa(C)$ is equivalent to the vanishing of its Abel–Jacobi image in the intermediate Jacobian. Using some formal considerations concerning isotrivial families of elliptic curves, we are able to bootstrap the calculations of [27] to determine the Abel-Jacobi image up to multiple and obtain the characterization of Theorem C.
In §3, we give two examples of genus 4 families of curves $\mathcal{C}\to B$ for which Theorem B applies. It would be interesting to completely characterize the vanishing of $\kappa(C)$ in these families as well, and to find more such families. The general proof strategy would follow that of Theorem C: using the Hodge conjecture, one shows that the vanishing of $\kappa(\mathcal{C}_{b})$ is equivalent to the vanishing of $\sigma(b)$, where $\sigma$ is a section of a certain abelian scheme $\mathcal{A}\to B$. (The analytification of $\sigma$ is the normal function associated to the Ceresa cycle over $B$, which lands in the ${\rm Aut}(C)$-fixed points of the intermediate Jacobian.) One then tries to identify $\mathcal{A}\to B$ and $\sigma$ explicitly.
@@ 1.3. Vanishing loci in genus $3$
Let $\mathcal{M}_{g}$ be the coarse moduli space of genus-$g$ curves in characteristic zero, seen as a variety over ${\rm\mathbb{Q}}$. Let $V_{g}^{\mathrm{rat}}\subset\mathcal{M}_{g}$ be the subset consisting of all (geometric) isomorphism classes of curves with vanishing $\kappa(C)$. Let $V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ be the analogous subset for $\bar{\kappa}(C)$. The sets $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ are both countable unions of closed algebraic subvarieties (Lemma 5.1). Moreover $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}$ and Ceresa’s result [14] shows that $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ for all $g\geq 3$. What can be said about the irreducible components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ and their dimensions? Collino and Pirola have shown that $V_{3}^{\mathrm{rat}}$ does not contain subvarieties of dimension $\geq 4$ that are not themselves contained in the hyperelliptic locus [16, Corollary 4.3.4]. Theorem C shows that $V_{3}^{\mathrm{alg}}$ contains the $2$-dimensional Picard locus and that $V_{3}^{\mathrm{rat}}$ contains infinitely many $1$-dimensional components.
@@
Here, $G_{48}$ is the group with GAP label $(48,33)$. The strata $X_{C_{9}}$ and $X_{G_{48}}$ are single closed points, represented by the curves $y^{3}=x^{4}-x$ and $y^{3}=x^{4}+1$ respectively.
Theorem D shows that $X_{G}\subset V_{3}^{\mathrm{rat}}$ if and only if $\operatorname{H}^{3}(J)^{G}=0$ for every (equivalently, some) curve in $X_{G}$, and that $X_{G}\subset V_{3}^{\mathrm{alg}}$ if and only if $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{G}=0$ for every curve in $X_{G}$. Thus, for $g=3$, the criteria of Theorems A and B exactly single out the strata $X_{G}$ where $\kappa(C)$ or $\bar{\kappa}(C)$ identically vanishes.
Finally, Shou-wu Zhang has recently announced a proof of a Northcott property for the Beilinson-Bloch height of the modified diagonal cycle $\Delta_{GKS}$ on $C^{3}$. More precisely, for each $g\geq 3$, there exists an open dense subset $U_{g}\subset\mathcal{M}_{g}$ such that for any $X\in{\rm\mathbb{R}}$ and $d\in{\rm\mathbb{N}}$, the number of $C\in U_{g}(\bar{\rm\mathbb{Q}})$ defined over a number field of degree at most $d$ and with $\langle\Delta_{GKS},\Delta_{GKS}\rangle<X$ is finite. In order to better understand Ceresa vanishing loci in families of curves, it would be of great interest to try to identify the largest such open dense set $U_{g}$, or equivalently, its complement $Z_{g}:=\mathcal{M}_{g}\setminus U_{g}$. The Northcott property implies that any positive dimensional component of $V_{g}^{\mathrm{rat}}$ is contained in $Z_{g}$, but there may be other components of $Z_{g}$ as well. Indeed, Theorem C shows that $X_{C_{3}}\subset Z_{3}$, even though $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$; see Remark 5.4.

### v2

@@ Abstract.
We prove cohomological vanishing criteria for the Ceresa cycle of a curve $C$ embedded in its Jacobian $J$. Namely:
@@ 1. $(A)$
    If $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$, then the Ceresa cycle is torsion modulo rational equivalence;
@@ 2. $(B)$
    If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then the Ceresa cycle is torsion modulo algebraic equivalence,
@@
We then use these criteria to study the simplest family of curves where $(B)$ holds but $(A)$ does not, namely the family of Picard curves $C\colon y^{3}=x^{4}+ax^{2}+bx+c$. Criterion $(B)$ and work of Schoen combine to show that the Ceresa cycle of a Picard curve is torsion in the Griffiths group. We furthermore determine exactly when it is torsion in the Chow group. As a byproduct, we deduce: there exist one-parameter families of plane quartic curves with torsion Ceresa Chow class; the torsion locus in $\mathcal{M}_{3}$ of the Ceresa Chow class contains infinitely many components; the order of a torsion Ceresa Chow class of a Picard curve over a number field $K$ is bounded, with the bound depending only on $[K\colon\mathbb{Q}]$. Finally, we determine which automorphism group strata are contained in the vanishing locus of the universal Ceresa cycle over $\mathcal{M}_{3}$.
@@ 1. Introduction
in the Chow group modulo rational equivalence. If $\kappa_{C,e}$ is torsion, then $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$, where $K_{C}$ is the canonical divisor class. Moreover if $(2g-2)e$ is canonical, then the image of $\kappa_{C,e}$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$ is independent of $e$ and we denote it by $\kappa(C)$, see §2.7 for these claims. Thus, the class $\kappa(C)$ vanishes if and only if $\kappa_{C,e}$ is torsion for some degree-$1$ divisor $e$.
@@
The vanishing of the Ceresa cycle is interesting for various reasons. For example, $\kappa(C)=0$ if and only if the Chow motive $\mathfrak{h}(C)$ has a multiplicative Chow-Künneth decomposition (by [22, Proposition 3.1] and Proposition 2.12). Moreover $\bar{\kappa}(C)=0$ if and only if the tautological subring modulo algebraic equivalence is generated by a theta divisor (by [6, Corollary 3.4]), in which case Poincaré’s formula $[C]=\frac{\Theta^{g-1}}{(g-1)!}$ holds modulo algebraic equivalence. More generally, the Ceresa cycle over $\mathcal{M}_{g}$ serves as a testing ground for the study of homologically trivial algebraic cycles in codimension greater than 1.
@@ 1.1. Vanishing criteria
This improves on a recent result of Qiu and Zhang stating that if $(\operatorname{H}^{1}(C)^{\otimes 3})^{{\rm Aut}(C)}=0$, then $\kappa(C)=0$ [43].111 This formulation is equivalent to theirs, using [51, Theorem 1.5.5]. By contrast, Theorem A requires only the weaker condition that the subrepresentation $\operatorname{H}^{3}(J)_{\mathrm{prim}}\subset\operatorname{H}^{3}(J)\simeq\bigwedge^{3}\operatorname{H}^{1}(C)\subset\operatorname{H}^{1}(C)^{\otimes 3}$ has no nontrivial ${\rm Aut}(C)$-fixed points. If the quotient $C/{\rm Aut}(C)$ has genus 00, then our hypothesis is equivalent to $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$.
@@
Our second result is a vanishing criterion for $\overline{\kappa}(C)$, however the result is conditional on the Hodge conjecture. In particular, we assume for the rest of the introduction that $k$ has characteristic 00.
@@
Assume the Hodge conjecture for abelian varieties. If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then $\bar{\kappa}(C)=0$.
More precisely, we require the Hodge conjecture for $J\times A$, where $A$ is the abelian variety described in Proposition 3.3. Note that $\operatorname{H}^{0}(J,\Omega_{J}^{3})\simeq\bigwedge^{3}\operatorname{H}^{0}(C,\Omega_{C}^{1})$, so the conditions of Theorems A and B both depend only on the abstract representation $(G,V)=({\rm Aut}(C),\operatorname{H}^{0}(C,\Omega_{C}^{1}))$.
@@ 1.2. Picard curves
The condition $\operatorname{H}^{3}(J)_{\mathrm{prim}}^{{\rm Aut}(C)}=0$ in Theorem A is only rarely satisfied, e.g. it holds for exactly two plane quartic curves over ${\rm\mathbb{C}}$ (see Theorem E below). The condition $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$ of Theorem B is satisfied more often; e.g. it holds for all Picard curves $y^{3}=x^{4}+ax^{2}+bx+c$. For such curves, it turns out that we require a nontrivial case of the Hodge conjecture to apply Theorem B, namely that a certain Weil class on an abelian fourfold is algebraic. Fortunately, Schoen has proved the Hodge conjecture in our specific situation [46]. (See also recent work of Markman [36].) This leads to an unconditional proof of the vanishing of $\bar{\kappa}(C)$. By further analyzing $\kappa(C)$, we prove the following precise characterization of the vanishing of $\kappa(C)$ in the Picard family:
@@
Let $C_{f}$ be a smooth projective Picard curve with model
@@
Here, ${\rm disc}(f)$ is the discriminant of $f$ and the coordinates of $P_{f}$ are the usual $I$- and $J$-invariants of the binary quartic form $z^{4}f(x/z)$; see §4.1. The fact that $P_{f}$ defines a point on $E_{f}$ follows from a classical relation between the invariants of a binary quartic. Picard curves have a unique point $\infty$ at infinity, for which $(2g-2)\infty=4\infty$ is canonical. For a more precise version of Theorem C relating the torsion orders of $\kappa_{C_{f},\infty}$ and $P_{f}$, see §4.2.
Theorem C generalizes the results of [31] concerning bielliptic Picard curves, i.e. those with $b=0$ in (1.2). There we exploited the explicit geometry of the bielliptic cover to characterize the vanishing locus of $\kappa(C)$. Here we must proceed more indirectly: using Schoen’s work on the Hodge conjecture, we first show that the vanishing of $\kappa(C)$ is equivalent to the vanishing of its Abel–Jacobi image in the intermediate Jacobian. Using the Shioda-Tate formula and considerations of isotrivial families of elliptic curves, we are able to bootstrap the calculations of [31] to determine the Abel-Jacobi image up to multiple and obtain the characterization of Theorem C.
In §3, we give several examples of higher genus families of curves $\mathcal{C}\rightarrow B$ for which Theorem B applies. It would be interesting to completely characterize the vanishing of $\kappa(C)$ in these families as well, and to find more such families. The general proof strategy would follow that of Theorem C: using the Hodge conjecture, one shows that the vanishing of $\kappa(\mathcal{C}_{b})$ is equivalent to the vanishing of $\sigma(b)$, where $\sigma$ is a section of a certain abelian scheme $\mathcal{A}\rightarrow B$. (The analytification of $\sigma$ is the normal function associated to the Ceresa cycle over $B$, which lands in the ${\rm Aut}(C)$-fixed points of the intermediate Jacobian.) One then tries to identify $\mathcal{A}\rightarrow B$ and $\sigma$ explicitly.
@@ 1.3. Vanishing loci in genus $3$
Let $\mathcal{M}_{g}$ be the coarse moduli space of genus-$g$ curves $C$, seen as a variety over ${\rm\mathbb{Q}}$. Let $V_{g}^{\mathrm{rat}}\subset\mathcal{M}_{g}$ be the locus of curves with $\kappa(C)=0$. Let $V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ be the analogous subset for $\bar{\kappa}(C)$. The subsets $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ are each a countable union of closed algebraic subvarieties (Lemma 5.1). We have $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}$, and Ceresa’s result [14] shows that $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ for all $g\geq 3$. What can be said about the irreducible components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ and their dimensions? For a fixed $g$, are there only finitely many components? Collino and Pirola showed that $V_{3}^{\mathrm{rat}}$ does not contain subvarieties of dimension $\geq 4$ that are not themselves contained in the hyperelliptic locus [16, Corollary 4.3.4]. On the other hand, Theorem C shows that $V_{3}^{\mathrm{alg}}$ contains the $2$-dimensional Picard locus. Turning to $V_{3}^{\mathrm{rat}}$, we answer several open question about its geometry and arithmetic by analyzing the torsion locus of the section $P_{f}$ of $E_{f}$:
Theorem D.
Let $X\subset\mathcal{M}_{3}$ be the Picard locus.
  1. $(i)$
    $(X\cap V_{3}^{\mathrm{rat}})\times_{{\rm\mathbb{Q}}}\mathbb{C}$ is a countably infinite disjoint union of rational curves. In particular, there exist one-parameter families of plane quartic curves in $V_{3}^{\mathrm{rat}}$, and $V_{3}^{\mathrm{rat}}$ contains infinitely many positive-dimensional components.
  2. $(ii)$
    If $K$ is a number field and $C\in(X\cap V_{3}^{\mathrm{rat}})(K)$, then the order of $\kappa_{C,\infty}$ in ${\rm CH}_{1}(J)$ is bounded above by a quantity depending only on $[K\colon{\rm\mathbb{Q}}]$.
An example of a one-parameter family as in $(i)$ is the family $C_{t}\colon y^{3}=x^{4}-12x^{2}+tx-12$. Note that we have previously shown in [31, Theorem 1.3] that the torsion orders of the classes $\kappa_{C,e}$, where $C$ has genus $3$ and $4e=K_{C}$, are unbounded.
@@ 2. $(ii)$
Here, $G_{48}$ is the group with GAP label $(48,33)$. The strata $X_{C_{9}}$ and $X_{G_{48}}$ are single closed points, represented by the curves $y^{3}=x^{4}+x$ and $y^{3}=x^{4}+1$ respectively.
Theorem E shows that $X_{G}\subset V_{3}^{\mathrm{rat}}$ if and only if $\operatorname{H}^{3}(J)^{G}=0$ for every (equivalently, some) curve in $X_{G}$, and that $X_{G}\subset V_{3}^{\mathrm{alg}}$ if and only if $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{G}=0$ for every curve in $X_{G}$. Thus, for $g=3$, the criteria of Theorems A and B exactly single out the strata $X_{G}$ where $\kappa(C)$ or $\bar{\kappa}(C)$ identically vanishes. Does this continue to hold in higher genus? Are there non-hyperelliptic curves of arbitrarily large genus for which Theorem A or B applies?
We also use Theorem B to find the first examples of positive dimensional components of $V_{g}^{\mathrm{alg}}$ with generic point satisfying ${\rm Aut}(C)=\{1\}$. [^2] Qiu and Zhang have previously found individual curves with $\kappa(C)=0$ and ${\rm Aut}(C)=\{1\}$ [44]. This uses the following fact: if $C$ dominates another curve $D$, then $\overline{\kappa}(C)=0$ implies $\overline{\kappa}(D)=0$. For example, the genus 6 family
$C_{t}\colon y^{9}=\left(\frac{x+1}{x-1}\right)\left(\frac{x+t}{x-t}\right)^{3},$
admits a $D_{9}$-action and lies in $V_{6}^{\mathrm{alg}}$ (conditional on the Hodge conjecture). The quotient by the involution $(x,y)\mapsto(-x,y^{-1})$ is the genus 3 family
$X_{u}\colon y^{4}=xy^{3}+ux^{3}+3y^{2}-2ux-uy-1,$
where $u=(t-1)/(t+1)$, which lies in $V_{3}^{\mathrm{alg}}$ by the fact above, and satisfies ${\rm Aut}(X_{u})=\{1\}$ generically. A similar construction gives a 1-dimensional component of $V_{4}^{\mathrm{alg}}$ (again conditional on the Hodge conjecture) with trivial generic automorphism group; see Theorem 3.11. Thus, even in genus 3, the implications of Theorems A and B are not yet fully understood, since it is hard to know which curves admit high genus covers that satisfy our vanishing criteria. We should remark that, as far as we know, there are no known examples of curves $C$ with $\overline{\kappa}(C)=0$, ${\rm Aut}(C)=\{1\}$, and ${\rm End}(J)={\rm\mathbb{Z}}$; see [20] for data confirming this for many quartic plane curves. The curves $X_{u}$ are no exceptions as generically we have ${\rm End}^{0}(\operatorname{Jac}(X_{u}))={\rm\mathbb{Q}}(\zeta_{9}+\zeta_{9}^{-1})$ by Remark 3.12.
Finally, Gao and Zhang have recently proven a Northcott property for the Beilinson-Bloch height of the Ceresa cycle on $J$ and (equivalently) the modified diagonal cycle $\Delta(C)$ on $C^{3}$ [24]. More precisely, for each $g\geq 3$, there exists an open dense subset $U_{g}\subset\mathcal{M}_{g}$ such that for any $X\in{\rm\mathbb{R}}$ and $d\in{\rm\mathbb{N}}$, the number of $C\in U_{g}(\bar{\rm\mathbb{Q}})$ defined over a number field of degree at most $d$ and with $\langle\Delta(C),\Delta(C)\rangle<X$ is finite. In order to better understand Ceresa vanishing loci in families of curves, it is of great interest to try to identify the largest such open dense set $U_{g}$, or equivalently, its complement $Z_{g}^{\mathrm{slim}}:=\mathcal{M}_{g}\setminus U_{g}$. The Northcott property implies that any positive dimensional component of $V_{g}^{\mathrm{rat}}$ is contained in $Z_{g}^{\mathrm{slim}}$, but $Z_{g}^{\mathrm{slim}}$ may be strictly larger than the union of the positive dimensional components of $V_{g}^{\mathrm{rat}}$. Indeed, Theorem D shows that $X_{C_{3}}\subset Z_{3}^{\mathrm{slim}}$, even though $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$.